BORWEIN AND BRADLEY’S APÉRY-LIKE

FORMULAE FOR ζ(4n + 3)

GERT ALMKVIST AND ANDREW GRANVILLE Abstract. We prove a formula for ζ(4n + 3) discovered by Borwein and Bradley (Experimental Mathematics 6:3 (1997), 181– 194).